You glance at the clock when you hear the first rumbling of an earthquake.

You feel the strong shaking exactly 13 seconds later. Eimate the likely distance of the quake from you.

Average velocity
P wave 8 km s-1
S wave 4 km s-1
Answer #1

d = (tS-tP/(1/vS-1/vP) , where tsand tp secondary and primary wve time

ts-tp= 13s , vs=4km/s, vp=8km/s

d=13/(1/4 -1/8) =104km